# What Can String Probability Tell Us About Grammaticality?

### v1

@@ 1 Introduction
Understanding what probabilistic language models (LMs) can learn about grammar has major ramifications for theories of language learning and structure (Linzen, 2019; Warstadt and Bowman, 2022; Baroni, 2022; Piantadosi, 2024). In the past decade, there have been many efforts to evaluate LMs’ grammatical knowledge (e.g., Warstadt et al., 2020; Hu et al., 2020; Linzen et al., 2016; Tjuatja et al., 2025), with some asserting that models have largely achieved grammatical competence (e.g., Mahowald et al., 2024) and others much more skeptical (e.g., Dentella et al., 2023; Lan et al., 2024; Fox and Katzir, 2024).
@@
Part of the field’s uncertainty over LMs’ grammatical competence stems from uncertainty over how to best assess grammatical knowledge in models. Given the success and convenience of prompting methods, a tempting approach is to simply “ask” models what sentences are grammatical or not (Dentella et al., 2023; Katzir, 2023), just as is commonly done for humans (Schütze, 2016; Sprouse and Almeida, 2012; Mahowald et al., 2016). But answering “Is this sentence grammatical?” requires more than just knowledge of grammar: it requires knowing what grammaticality means, as well as other auxiliary abilities such as being able to (truthfully) answer questions. As a result, this method systematically underestimates grammatical competence in LMs (Hu and Levy, 2023; Hu et al., 2024). It’s easy to see the problem if we imagine a model trained without ever seeing the word “grammatical”: it would have the same underlying knowledge of linguistic structure but be unable to answer the question.
@@
An alternate approach is to measure the probabilities that models assign to strings, with the logic that models should assign higher probability to grammatical versus ungrammatical strings. But it is not immediately obvious that this is the best way to assess LMs’ grammatical abilities, as grammaticality and probability are fundamentally distinct notions in linguistics (Chomsky, 1957; Berwick, 2018). A well-known illustration of this distinction is the sentence “Colorless green ideas sleep furiously” (Chomsky, 1957). The string has low probability (at least when it was originally coined), but, crucially, people still have the intuition that it is grammatical in a way “*Furiously sleep ideas green colorless” isn’t. This line of thinking might make it seem like assessing models via string probability is fundamentally flawed. Indeed, critics have argued that the distinction between likelihood and grammaticality is “entirely foreign” (Katzir, 2023) to LMs, making them unsuitable models of grammatical competence.
@@
However, Fox and Katzir (2024), Lan et al. (2024), and others go on to note that, in some cases, probability may be aligned enough with grammatically that it can be informative. And in practice, probability-based evaluations of grammatical knowledge in LMs use minimal pairs—i.e., pairs of sentences that differ only slightly from each other, and which form a grammaticality contrast (Marvin and Linzen, 2018; Futrell et al., 2019; Warstadt et al., 2020; Hu et al., 2020; Wilcox et al., 2023; Hu et al., 2024). Intuitively, researchers construct minimal pairs to isolate a specific grammatical contrast and factor out other properties that might affect string probability, such as length or lexical frequencies. But even this practice has also been criticized by recent work. For example, Leivada et al. (2024a) write: “If LMs need specific comparisons in order to tell apart grammatical from ungrammatical sentences, this already counts as an inherent discrepancy from humans, who are able to make such judgments without such a comparison”. If string probability offers a window into grammaticality, they argue, then it should be possible to find a threshold on probability that separates grammatical and ungrammatical sentences (Leivada et al., 2024a, b).
@@
In the rest of this paper, we first give a formal characterization of string probabilities in corpora and models. Then, we derive three predictions and test them empirically on 280K sentence pairs in English and Chinese. First, we predict a correlation between the log-probability of grammatical and ungrammatical sentences that convey roughly the same message, since their message probability is controlled for. Second, we predict that differences in human acceptability judgments on appropriately controlled minimal pairs will align with differences in log-probability of the strings in the pair. Finally, we predict a lack of separation in string probability space between unpaired grammatical and ungrammatical sentences. While this phenomenon has been taken to indicate a failure of models to capture grammaticality (Leivada et al., 2024b, a), we argue that it follows from our framework under reasonable assumptions.
More generally, we see our contribution as providing theoretical grounding for the practice of using minimal-pair probability comparisons to assess LMs’ grammatical knowledge. While this practice is widely used in NLP, it has generally been given only brief and informal justification in empirical work (e.g., Warstadt et al., 2020), or rejected altogether (e.g., Leivada et al., 2024a). We use our theoretical analysis and empirical results to motivate recommendations for future work on grammatical evaluation of LMs.
@@ 2.1.1 Toy example
$\;\;\;\;\;\;\;\;\;\;\;\;\land\text{moon}(y)\land\text{emerge}(y)]$
@@
Although for a given m{m} these strings are all errorful realizations, they might not all be equally likely. In our framework, string probability is influenced by the likelihood of (potential) underlying messages. Here, we expect $P({m}_{1})>P({m}_{2})>P({m}_{3})$: i.e., it is most probable (at least on Earth) to express a message about a unique moon, somewhat probable to express a message about a single moon, and improbable to express a message about multiple moons. These differences in message probabilities can conflict with grammaticality in practice. LABEL:fig:gpt2-example shows surprisals (i.e., negative log probability) assigned by GPT-2 (Radford et al., 2019) to each of the 8 toy strings from LABEL:fig:sentence-grid. While the 3 grammatical strings have lower surprisals on average than the 5 ungrammatical strings, we also see that the strings which mention a singular moon tend to have lower surprisal. For example, GPT-2 assigns higher surprisal to the grammatical “The moons emerge” than the ungrammatical “*The moon emerge”.
@@ 2.2 Error model
The role of errors in speech comprehension and production has been studied extensively (e.g., Levy, 2008; Goldrick, 2011). We adopt a set of minimal working assumptions required for a basic error model. Let ${\mathbf{s}}^{*({m})}$ be the “grammatical realization” of $m$: i.e., the string $\mathbf{s}$ such that $P({\mathbf{s}}|{m},{G}=1)=1$. By ˜1, this grammatical realization is unique. Then, $P({\mathbf{s}}|{m},{G}=0)$ is concentrated in an “error neighborhood” of ${\mathbf{s}}^{*({m})}$ that excludes ${\mathbf{s}}^{*({m})}$. That is, although violating a language’s syntax, an ungrammatical realization of a message $m$ tends to be mostly similar to a grammatical realization of $m$.
@@ 3 Three predictions
Native speaker acceptability judgments vary with both grammatical well-formedness and meaning plausibility (Schütze, 2016). Using our framework, we operationalize (i) with $\log P({m})$, and (ii) with the number of errors. Then if we consider minimal pairs $({\mathbf{s}},{\mathbf{s}}^{\prime})$, where the understood message between $\mathbf{s}$ and $\mathbf{s}$$^{\prime}$ is the same, the difference in acceptability judgments depends primarily on the error probability of taking $\mathbf{s}$ to $\mathbf{s}$$^{\prime}$, as does the difference in string log-probability. We therefore predict that differences in string probability are correlated with differences in human acceptability judgments, within minimal pairs (˜2a). As the “minimalness” of the pair decreases, the contribution of message M{M} increases, and we expect to find weaker correlation between log-probability differences and acceptability judgment differences (˜2b).
@@
BLiMP English 66993 Warstadt et al. (2020)
SCaMP-P English 67000 McCoy and Griffiths (2025)
SCaMP-I English 67000 McCoy and Griffiths (2025)
SyntaxGym English 1018 Hu et al. (2020)
ZhoBLiMP Chinese 35400 Liu et al. (2024)
SLING Chinese 39976 Song et al. (2022)
CoLA English 8551 Warstadt et al. (2019)
LI English 1883 Sprouse et al. (2013); Mahowald et al. (2016)
HLL Chinese 213 Chen et al. (2020)
@@
In illustrating the distinction between probability and grammaticality, Chomsky (1957) wrote: “If we rank the sequences of a given length in order of statistical approximation to English, we will find both grammatical and ungrammatical sequences scattered throughout the list”. By viewing probability as influenced by grammaticality and meaning, our framework provides a theoretical basis for Chomsky’s prediction: i.e., string probability does not separate grammatical and ungrammatical strings. While Leivada et al. (2024a, b) argue that this lack of separation means that it is problematic to use probability to measure grammaticality, we note that this prediction falls directly out of our theoretical framework.
@@
We note that so far we have only discussed pure string probability, which has been investigated in recent studies (Leivada et al., 2024a, b). There are reasons to expect that grammatical/ungrammatical strings would not be separated by pure probability: when there is no upper bound on the length of grammatical strings and at least some ungrammatical string is assigned non-zero probability (as is the case for LMs), then there must be an infinite set of grammatical strings that are assigned lower probability than that ungrammatical string. Indeed, our framework also suggests that replacing pure string probability with a “normalizing” function of string probability and other grammar-independent features should bring the distributions of messages associated with grammatical and ungrammatical strings closer together, and thereby increase the separability. This prediction provides a novel justification for previous work which has hypothesized that grammaticality and probability are linked through a complex function (Pauls and Klein, 2012; Lau et al., 2017; Tjuatja et al., 2025).
@@ 4.1 Evaluation materials
To test ˜1, we need a set of paired grammatical and ungrammatical sentences with varying degrees of “minimalness”. We test our theory on existing data sets from two typologically different languages, English and Mandarin Chinese, as our theoretical framework is language-agnostic and only makes basic assumptions about the generative process of corpus strings. We evaluate models on five datasets, summarized in LABEL:tab:datasets: BLiMP (Warstadt et al., 2020), SCaMP (McCoy and Griffiths, 2025), and SyntaxGym [^3] We use the subset of SyntaxGym compatible with sentence scoring. (Hu et al., 2020) in English, and ZhoBLiMP (Liu et al., 2024) and SLING (Song et al., 2022) in Chinese. Each dataset is proposed to contain “minimal pairs”, although the pairs may diverge to varying degrees from the theoretical ideal defined in Definition˜4.
@@ 4.3 Computing string probability
The predictions of our framework apply to any probabilistic model that has learned a reasonably accurate distribution of $P(S)$. We felt this to be a reasonable assumption for moderately-sized Transformer models trained on Internet-scale text. Although an interesting avenue for further research, directly testing the influence of specific factors such as model size was not the key motivation of our experiments, so we simply chose two models for each language, covering multiple sizes and model families. We evaluated two open-source base (i.e., not fine-tuned) models of varying sizes on each dataset (see LABEL:tab:models for details). We evaluated GPT-2 (Radford et al., 2019) and Llama-3-70B (AI@Meta, 2024) on the English datasets. For the Chinese datasets, we evaluated a GPT-2 model trained on CLUECorpusSmall (Xu et al., 2020), which we refer to as GPT-2 ZH (Zhao et al., 2019), and Llama-3-8B ZH (Cui et al., 2023).
@@ 5.1 Evaluation materials
To test ˜2, we need human acceptability judgments of each sentence in isolation. Our evaluation datasets are summarized in LABEL:tab:datasets. We only used existing data and did not collect any new human judgments. The English dataset (LI) includes pairs from Sprouse et al. (2013) and Mahowald et al. (2016). Both studies randomly sampled paired sentences (grammatical vs. ungrammatical/questionable) from Linguistic Inquiry journal papers and collected acceptability judgments for each sentence from native English speakers. The Chinese dataset (HLL) includes pairs from Chen et al. (2020), where the authors collected acceptability judgments for each sentence within related groups (pairs, triples, or $n$-tuples) from a Chinese syntax textbook (Huang et al., 2009). [^5] For the groups with more than two sentences, we created pairs by juxtaposing all possible grammatical vs. ungrammatical/questionable pairings. That is, for a given group with grammatical sentences $\{g_{1},\dots,g_{n}\}$ and ungrammatical sentences $\{u_{1},\dots,u_{m}\}$ we create all the possible pairs $(g_{i},u_{j})$ for $i\in[1,n]$ and $j\in[1,m]$; usually there are 2-4 sentences in a group.
@@ 5.2 Results
The results for these analyses are displayed in Figure˜4. LABEL:fig:model-human-alignment-scatter shows that human Likert score differences are correlated with log probability differences (˜2). We also note that similar results have been reported for other languages (Suijkerbuijk et al., 2025).
@@ 6 Prediction 3: Poor separation of grammatical/ungrammatical strings
We now test our final prediction: potentially poor separation between the probability of grammatical and ungrammatical strings. As discussed in Section˜3 and Section˜A.3, our framework predicts that the limiting factor for such separation is the variance of messages associated with grammatical and ungrammatical string sets. Therefore, in addition to examining raw string probability, in this section we introduce several normalizing transformations that reduce variance in messages and should thereby also increase separability. Specifically, we propose a novel scoring function that represents the Bayes factor between different generative processes of observed strings. And we give a novel derivation for the Syntactic Log Odds Ratio (SLOR; Pauls and Klein, 2012; Lau et al., 2017), as equivalent to the average Pointwise Mutual Information between a word and its preceding context. We find that neither raw string probability nor any normalized string probabilities result in good separation of grammatical/ungrammatical strings, in line with our prediction.

### v2

@@ 1 Introduction
Understanding what probabilistic language models (LMs) can learn about grammar has major ramifications for theories of language learning and structure (linzen_what_2019; warstadt_what_2022; baroni_proper_2022; piantadosi_modern_2023). In the past decade, there have been many efforts to evaluate LMs’ grammatical knowledge (e.g., warstadt_blimp_2020; hu_systematic_2020; linzen_assessing_2016; tjuatja_what_2024), with some asserting that models have largely achieved grammatical competence (e.g., mahowald_dissociating_2024) and others much more skeptical (e.g., dentella_systematic_2023; lan_large_2024; fox2024large).
@@
Part of the field’s uncertainty over LMs’ grammatical competence stems from uncertainty over how to best assess grammatical knowledge in models. Given the success and convenience of prompting methods, a tempting approach is to simply “ask” models what sentences are grammatical or not (dentella_systematic_2023; katzir_why_2023), just as is commonly done for humans (schutze_empirical_2016; sprouse2012assessing; mahowald2016snap). But answering “Is this sentence grammatical?” requires more than just knowledge of grammar: it requires knowing what grammaticality means, as well as other auxiliary abilities such as being able to (truthfully) answer questions. As a result, this method systematically underestimates grammatical competence in LMs (hu_prompting_2023; hu_language_2024). It’s easy to see the problem if we imagine a model trained without ever seeing the word “grammatical”: it would have the same underlying knowledge of linguistic structure but be unable to answer the question.
@@
An alternate approach is to measure the probabilities that models assign to strings, with the logic that models should assign higher probability to grammatical versus ungrammatical strings. But it is not immediately obvious that this is the best way to assess LMs’ grammatical abilities, as grammaticality and probability are fundamentally distinct notions in linguistics (chomsky_syntactic_1957; berwick_syntactic_2018). A well-known illustration of this distinction is the sentence “Colorless green ideas sleep furiously” (chomsky_syntactic_1957). The string has low probability (at least when it was originally coined), but, crucially, people still have the intuition that it is grammatical in a way “*Furiously sleep ideas green colorless” isn’t. This line of thinking might make it seem like assessing models via string probability is fundamentally flawed. Indeed, critics have argued that the distinction between likelihood and grammaticality is “entirely foreign” (katzir_why_2023) to LMs, making them unsuitable models of grammatical competence.
@@
However, fox2024large, lan_large_2024, and others go on to note that, in some cases, probability may be aligned enough with grammatically that it can be informative. And in practice, probability-based evaluations of grammatical knowledge in LMs use minimal pairs—i.e., pairs of sentences that differ only slightly from each other, and which form a grammaticality contrast (marvin_targeted_2018; futrell_neural_2019; warstadt_blimp_2020; hu_systematic_2020; wilcox_using_2023; hu_language_2024). Intuitively, researchers construct minimal pairs to isolate a specific grammatical contrast and factor out other properties that might affect string probability, such as length or lexical frequencies. But even this practice has also been criticized by recent work. For example, leivada_evaluating_2024 write: “If LMs need specific comparisons in order to tell apart grammatical from ungrammatical sentences, this already counts as an inherent discrepancy from humans, who are able to make such judgments without such a comparison”. If string probability offers a window into grammaticality, they argue, then it should be possible to find a threshold on probability that separates grammatical and ungrammatical sentences (leivada_evaluating_2024; leivada_reply_2024).
@@
In the rest of this paper, we first give a formal characterization of string probabilities in corpora and models. Then, we derive three predictions and test them empirically on 280K sentence pairs in English and Chinese. First, we predict a correlation between the log-probability of grammatical and ungrammatical sentences that convey roughly the same message, since their message probability is controlled for. Second, we predict that differences in human acceptability judgments on appropriately controlled minimal pairs will align with differences in log-probability of the strings in the pair. Finally, we predict a lack of separation in string probability space between unpaired grammatical and ungrammatical sentences. While this phenomenon has been taken to indicate a failure of models to capture grammaticality (leivada_reply_2024; leivada_evaluating_2024), we argue that it follows from our framework under reasonable assumptions.
More generally, we see our contribution as providing theoretical grounding for the practice of using minimal-pair probability comparisons to assess LMs’ grammatical knowledge. While this practice is widely used in NLP, it has generally been given only brief and informal justification in empirical work (e.g., warstadt_blimp_2020), or rejected altogether (e.g., leivada_evaluating_2024). We use our theoretical analysis and empirical results to motivate recommendations for future work on grammatical evaluation of LMs.
@@ 2.1.1 Toy example
$\;\;\;\;\;\;\;\;\;\;\;\;\land\text{moon}(y)\land\text{emerge}(y)$
$\;\;\;\;\;\;\;\;\;\;\;\;\land(x\neq y)]$
@@
Although for a given m{m} these strings are all errorful realizations, they might not all be equally likely. In our framework, string probability is influenced by the likelihood of (potential) underlying messages. Here, we expect $P({m}_{1})>P({m}_{2})>P({m}_{3})$: i.e., it is most probable (at least on Earth) to express a message about a unique moon, somewhat probable to express a message about a single moon, and improbable to express a message about multiple moons. These differences in message probabilities can conflict with grammaticality in practice. LABEL:fig:gpt2-example shows surprisals (i.e., negative log probability) assigned by GPT-2 (radford_language_2019) to each of the 8 toy strings from LABEL:fig:sentence-grid. While the 3 grammatical strings have lower surprisals on average than the 5 ungrammatical strings, we also see that the strings which mention a singular moon tend to have lower surprisal. For example, GPT-2 assigns higher surprisal to the grammatical “The moons emerge” than the ungrammatical “*The moon emerge”.
@@ 2.2 Error model
The role of errors in speech comprehension and production has been studied extensively (e.g., levy-2008-noisy; goldrick2011errors). We adopt a set of minimal working assumptions required for a basic error model. Let ${\mathbf{s}}^{*({m})}$ be the “grammatical realization” of $m$: i.e., the string $\mathbf{s}$ such that $P({\mathbf{s}}|{m},{G}=1)=1$. By ˜1, this grammatical realization is unique. Then, $P({\mathbf{s}}|{m},{G}=0)$ is concentrated in an “error neighborhood” of ${\mathbf{s}}^{*({m})}$ that excludes ${\mathbf{s}}^{*({m})}$. That is, although violating a language’s syntax, an ungrammatical realization of a message $m$ tends to be mostly similar to a grammatical realization of $m$.
@@ 3 Three predictions
Native speaker acceptability judgments vary with both grammatical well-formedness and meaning plausibility (schutze_empirical_2016). Using our framework, we operationalize (i) with $\log P({m})$, and (ii) with the number of errors. Then if we consider minimal pairs $({\mathbf{s}},{\mathbf{s}}^{\prime})$, where the understood message between $\mathbf{s}$ and $\mathbf{s}$$^{\prime}$ is the same, the difference in acceptability judgments depends primarily on the error probability of taking $\mathbf{s}$ to $\mathbf{s}$$^{\prime}$, as does the difference in string log-probability. We therefore predict that differences in string probability are correlated with differences in human acceptability judgments, within minimal pairs (˜2a). As the “minimalness” of the pair decreases, the contribution of message M{M} increases, and we expect to find weaker correlation between log-probability differences and acceptability judgment differences (˜2b).
@@
BLiMP English 66993 warstadt_blimp_2020
SCaMP-P English 67000 mccoy_modeling_2025
SCaMP-I English 67000 mccoy_modeling_2025
SyntaxGym English 1018 hu_systematic_2020
ZhoBLiMP Chinese 35400 liu_zhoblimp_2024
SLING Chinese 39976 song_sling_2022
CoLA English 8551 warstadt_neural_2019
LI English 1883 sprouse2013comparison; mahowald2016snap
HLL Chinese 213 chen2020assessing
@@
In illustrating the distinction between probability and grammaticality, chomsky_syntactic_1957 wrote: “If we rank the sequences of a given length in order of statistical approximation to English, we will find both grammatical and ungrammatical sequences scattered throughout the list”. By viewing probability as influenced by grammaticality and meaning, our framework provides a theoretical basis for Chomsky’s prediction: i.e., string probability does not separate grammatical and ungrammatical strings. While leivada_evaluating_2024; leivada_reply_2024 argue that this lack of separation means that it is problematic to use probability to measure grammaticality, we note that this prediction falls directly out of our theoretical framework.
@@
We note that so far we have only discussed pure string probability, which has been investigated in recent studies (leivada_evaluating_2024; leivada_reply_2024). There are reasons to expect that grammatical/ungrammatical strings would not be separated by pure probability: when there is no upper bound on the length of grammatical strings and at least some ungrammatical string is assigned non-zero probability (as is the case for LMs), then there must be an infinite set of grammatical strings that are assigned lower probability than that ungrammatical string. Indeed, our framework also suggests that replacing pure string probability with a “normalizing” function of string probability and other grammar-independent features should bring the distributions of messages associated with grammatical and ungrammatical strings closer together, and thereby increase the separability. This prediction provides a novel justification for previous work which has hypothesized that grammaticality and probability are linked through a complex function (pauls_large-scale_2012; lau_grammaticality_2017; tjuatja_what_2024).
@@ 4.1 Evaluation materials
To test ˜1, we need a set of paired grammatical and ungrammatical sentences with varying degrees of “minimalness”. We test our theory on existing data sets from two typologically different languages, English and Mandarin Chinese, as our theoretical framework is language-agnostic and only makes basic assumptions about the generative process of corpus strings. We evaluate models on five datasets, summarized in LABEL:tab:datasets: BLiMP (warstadt_blimp_2020), SCaMP (mccoy_modeling_2025), and SyntaxGym [^3] We use the subset of SyntaxGym compatible with sentence scoring. (hu_systematic_2020) in English, and ZhoBLiMP (liu_zhoblimp_2024) and SLING (song_sling_2022) in Chinese. Each dataset is proposed to contain “minimal pairs”, although the pairs may diverge to varying degrees from the theoretical ideal defined in Definition˜4.
@@ 4.3 Computing string probability
The predictions of our framework apply to any probabilistic model that has learned a reasonably accurate distribution of $P(S)$. We felt this to be a reasonable assumption for moderately-sized Transformer models trained on Internet-scale text. Although an interesting avenue for further research, directly testing the influence of specific factors such as model size was not the key motivation of our experiments, so we simply chose two models for each language, covering multiple sizes and model families. We evaluated two open-source base (i.e., not fine-tuned) models of varying sizes on each dataset (see LABEL:tab:models for details). We evaluated GPT-2 (radford_language_2019) and Llama-3-70B (aimeta_llama_2024) on the English datasets. For the Chinese datasets, we evaluated a GPT-2 model trained on CLUECorpusSmall (xu_cluecorpus2020_2020), which we refer to as GPT-2 ZH (zhao2019uer), and Llama-3-8B ZH (cui_efficient_2023).
@@ 5.1 Evaluation materials
To test ˜2, we need human acceptability judgments of each sentence in isolation. Our evaluation datasets are summarized in LABEL:tab:datasets. We only used existing data and did not collect any new human judgments. The English dataset (LI) includes pairs from sprouse2013comparison and mahowald2016snap. Both studies randomly sampled paired sentences (grammatical vs. ungrammatical/questionable) from Linguistic Inquiry journal papers and collected acceptability judgments for each sentence from native English speakers. The Chinese dataset (HLL) includes pairs from chen2020assessing, where the authors collected acceptability judgments for each sentence within related groups (pairs, triples, or $n$-tuples) from a Chinese syntax textbook (huang2009syntax). [^5] For the groups with more than two sentences, we created pairs by juxtaposing all possible grammatical vs. ungrammatical/questionable pairings. That is, for a given group with grammatical sentences $\{g_{1},\dots,g_{n}\}$ and ungrammatical sentences $\{u_{1},\dots,u_{m}\}$ we create all the possible pairs $(g_{i},u_{j})$ for $i\in[1,n]$ and $j\in[1,m]$; usually there are 2-4 sentences in a group.
@@ 5.2 Results
The results for these analyses are displayed in Figure˜4. LABEL:fig:model-human-alignment-scatter shows that human Likert score differences are correlated with log probability differences (˜2). We also note that similar results have been reported for other languages (suijkerbuijk_blimp-nl_2024).
@@ 6 Prediction 3: Poor separation of grammatical/ungrammatical strings
We now test our final prediction: potentially poor separation between the probability of grammatical and ungrammatical strings. As discussed in Section˜3 and Section˜A.3, our framework predicts that the limiting factor for such separation is the variance of messages associated with grammatical and ungrammatical string sets. Therefore, in addition to examining raw string probability, in this section we introduce several normalizing transformations that reduce variance in messages and should thereby also increase separability. Specifically, we propose a novel scoring function that represents the Bayes factor between different generative processes of observed strings. And we give a novel derivation for the Syntactic Log Odds Ratio (SLOR; pauls_large-scale_2012; lau_grammaticality_2017), as equivalent to the average Pointwise Mutual Information between a word and its preceding context. We find that neither raw string probability nor any normalized string probabilities result in good separation of grammatical/ungrammatical strings, in line with our prediction.